# Kotko,I.,${}^{1}$ Banerjee,S.,${}^{2,3}$ Belczynski,K.,${}^{1}$

### v1

@@ 2.3.2 The impact of natal kicks on Gaia BH-like binaries
The natal kick velocity $v_{\mathrm{kick}}$ is the vector that plays an important if not the decisive role in the Gaia BH-like systems formation. What counts is not only the magnitude of the kick but also its direction and relative position of the stars on their mutual orbit when SN takes place.
@@
The natal kick of a given magnitude determines $v_{\mathrm{sys}}$ but not the eccentricity of the binary what is demonstrated in Fig. 5. However, there is a correlation between the eccentricity, post- and pre-SN orbital separation (see e.g. Kalogera, 1996) which is followed by our population of Gaia BH-like binaries as is shown in Fig. 6. For Gaia BH1, where the companion star is still on its main sequence, the evolution has not have enough time to change the binary orbit from its immediate post-SN configuration. The system may have widen or shrunken due to the natal kick but the total amount of the orbital separation change is tightly connected to the system post-SN eccentricity. If the natal kick causes binary orbit to widen the orbital separation may increase no more than twice ($a_{\mathrm{post-SN}}/a_{\mathrm{pre-SN}}\sim 2$) for binaries that have eccentricities $e\in[0.4;0.6]$. If the orbit shrinks due to the natal kick then it can shrink by no more than $40\%$ ($1.00.6<a_{\mathrm{post-SN}}/a_{\mathrm{pre-SN}}<$ 1.0) (see Fig. 6) for such binaries.
@@
Whether the binary shrinks or widens after SN depends on the components of the vector position of the exploding star on the orbit around its companion at the moment of SN explosion and on the components of natal kick vector, where the components of both vectors are drawn from distributions described in Sec. 2.1.
Apart from the shrinkage/widening of the orbit the natal kick changes the angle between the direction of BH spin and the direction of the binary angular momentum ($\theta$). We show the relation between angles $\theta$, $\gamma$ and $v_{\mathrm{sys}}$ (top) and $a_{\mathrm{pre-SN}}$/$a_{\mathrm{post-SN}}$ (bottom) for Gaia BH1-like and Gaia BH2-like binaries in Fig. 7. On both plots Gaia BH1-like binaries are marked with stars and Gaia BH2-like binaries are marked with filled circles. The formation of Gaia BH1-like binaries depends differently on $\gamma$ than the formation of Gaia BH2-like binaries and so does $\theta$.
@@
In the case of Gaia BH2-like binaries $\gamma$ is a major factor determining their population. Most of them ($\sim 95\%$) form if natal kick direction is close to the orbital plane ($\gamma\sim 90^{\circ}\pm 15$) and $v_{\mathrm{kick}}<41.6$ km/s. The binaries that survive SN and form Gaia BH2-like systems have $\theta$ in the whole range of values. It can be clearly seen from the $\gamma$ and $\theta$ distributions shown in Fig. 9: $\theta$ distribution is almost flat and $\gamma$ distribution is centred around $90^{\circ}$.
@@
2.3.3 Common envelope parameters $\lambda$ and $\alpha_{\mathrm{ce}}$
@@
The standard version of StarTrack has been using the $\lambda$ formulae from Xu & Li (2010a); Xu & Li (2010b). In course of work on this paper we found that for the stars with ZAMS masses in the range $\sim 35-45\,\mathrm{M_{\odot}}$, which are the progenitors of $\sim 8-10\,\mathrm{M_{\odot}}$ BHs, Xu & Li (2010b) prescription underestimated the binding energy stars at the core He-buring stage of their evolution, resulting in the high $\lambda$ values ($\lambda>4.0$). The formulae were fitted to the models calculated with the old Eddington’s code EV and we learnt (Xiang-Dong Li private communication) that they were updated using Modules for Experiments in Stellar Astrophysics code (MESA, Paxton et al. (2011)) in 2016 by Wang et al. (2016) and should be used in their most recent form. In calculations done for the purpose of this paper we incorporated the formulate from Wang et al. (2016) for stars more massive than $18.0\,\mathrm{M_{\odot}}$ (BH progenitors). We chose the formulae which were calculated using the prescription for the wind mass loss rates from Hurley et al. (2000) and Vink et al. (2001) (Wind1 in Wang et al. (2016)) and we take $\lambda$ as the average of $\lambda_{\mathrm{g}}$ (calculated taking into account gravitational binding energy only), $\lambda_{\mathrm{b}}$ (calculated taking into account the total energy, that is both gravitational and internal energy which consists of thermal energy, radiation energy, ionization and dissociation energy) and $\lambda_{\mathrm{h}}$ (which in addition to the total energy takes into account the enthalpy of the stellar envelope).
One should be aware that the results may vary depending which $\lambda$ prescription is used in the calculations. To see the evolution of stellar envelope binding energy for different $\lambda$ prescriptions we refer to Iorio et al. (2023) where envelope binding energy is plotted in Fig. A1 in Appendix. Unfortunately, it does not take into account the updated Wang et al. (2016) formulae.
@@
We would like to point out that our results are obtained with standard value of $\alpha_{\mathrm{ce}}=1.0$. There are analytical estimations done for the final orbital separation after common envelope phase in pre-Gaia BH1 systems, however they are based on the assumption that the star has to loose $20\,\mathrm{M_{\odot}}$ of mass in the envelope during CE which high overestimation when stellar winds are taken into account. The mass of BH-progenitor prior to CE stage is much lower: a star that starts on ZAMS with $\sim 40\,\mathrm{M_{\odot}}$ looses $\sim 50\%$ of its initial mass in the stellar winds by the time CE phase starts. Consequently, the envelope that has to be unbound during CE has mass of only $\sim 3\,\mathrm{M_{\odot}}$. If $M_{\mathrm{env}}=3.0\,\mathrm{M_{\odot}}$ is taken for calculations the post-CE orbital separation of the binary is $>100\,\mathrm{R_{\odot}}$ instead of a few $\mathrm{R_{\odot}}$. Invoking $\alpha_{\mathrm{ce}}>1$ is not necessary to form Gaia BH-like binaries in IBE channel.
@@ 2.3.4 Gaia BH1 as Low Mass X-ray Binary
Although there is no X-ray emission detected from Gaia BH1 and Gaia BH2 (Rodriguez et al., 2023), it was found that during the post-dormant BH binary evolution of both of those binaries they may become the symbiotic X-ray binaries with the subsequent phase of stable Roche-lobe overflow (RLOF) Rodriguez et al. (2023) becoming a a low mass X-ray binary (LMXB). The population of LMXBs that originated from Gaia BH1-like binaries would have long orbital periods ($P_{\mathrm{orb}}>100$ days).
@@ 3.2 Results: formation rate of field BH-star binaries
As evident from Figs. 10 and 11, the vast majority of dynamically assembled field BH-star systems have $M_{\rm comp}<3\mathrm{M_{\odot}}$. Nearly half of them have zero or small eccentricity, an MS companion, and $P_{\rm orb}<1$ day — these very tight binaries are outcomes of tidal capture between the BH and the companion star inside the parent cluster and the binary’s subsequent tidal circularization [^3] In the Ba21 runs, tidal energy loss and analytical tidal circularization were enabled (option $27=1$; Aarseth 2003).. Inside clusters, such close binaries behave like single, massive objects most of the time and are ejected dynamically in ways similar to those of ejection of single BHs (Kulkarni et al., 1993). These systems are potentially tidally interacting or mass transferring and are not compatible with the Gaia detected, detached BH-star binaries. There are also a few wide, $P_{\rm orb}\gtrsim 100$ day, MS-companion binaries — these systems are assembled in the cluster via dynamical exchange encounters and are also ejected via dynamical scatterings (see above), making them eccentric in general.
@@ 4 Conclusions
In this paper we have investigated two possible channels of formation of two dormant BH binaries Gaia BH1 and Gaia BH2: the isolated binary evolution (Sec. 2) and the dynamical interactions in young massive and open star clusters (Sec. 3). We used the population synthesis code StarTrack for IBE calculations (Sec. 2.1) and updated star-by-star N-body integration code NBODY7 for cluster models (Sec. 3.1). The two systems from IBE Gaia-BH like binaries population, that match well the parameters of Gaia BH1 and Gaia BH2, were evolved till the Hubble time to find the characteristics of low-mass X-ray binaries that possibly form from such systems (Sec.2.3.4). We have also analysed what are the constrains on magnitude and direction of the natal kick imparted to BH set by the orbital parameters and systemic velocities of Gaia BH1-like and Gaia BH2-like binaries (Sec. 2.2.3 and Sec. 2.3.2).
@@
The formation rate of Gaia BH-like binaries inferred from the simulations is sensitive to the criteria adopted to classify a binary as Gaia BH-like (Sec. 2.2.1). In IBE channel it can differ by three orders of magnitude depending on what we consider a Gaia BH-like binary (see Table 3). Ensuring that the common criteria are set and the same total stellar mass per simulation is adopted for IBE and dynamical interactions in young clusters, we find that the formation rate of Gaia BH-like binaries for both channels is equal $2.7\times 10^{-7}\,\mathrm{M_{\odot}}^{-1}$. If we consider the general population of field BH-low mass star ($<3\,\mathrm{M_{\odot}}$) binaries from young massive and open clusters and corresponding BH-low mass star population from IBE the formation rate of such binaries in both channels is on order of $\sim 10^{-6}\,\mathrm{M_{\odot}}^{-1}$.
@@
  • The population of Gaia BH1-like binaries is not vulnerable to the natal kick velocity distribution adopted but a sevenfold increase in the number of Gaia BH2-like binaries in model V2 comparing to model V1 favours the two-component Maxwellian distribution with low- and high-velocity component ($\sigma\approx 21$ km/s and $\sigma\approx 107$ km/s) suggested by Zhao et al. (2023).
  • The median value of natal kick velocity imparted to the systems that become Gaia BH1-like and Gaia BH2-like binaries is $\sim 39$ km/s for the former and $\sim 19$ km/s for the latter (assuming the natal kick distribution from Zhao et al. (2023)).
@@
  • For Gaia BH1-like binary to be formed the kicks should be directed out of the binary orbital plane: the higher is the natal kick velocity the further from orbital plane should the kick be directed.

### v2

@@ 2.3.3 The evolution Gaia BH-like binaries
The natal kick velocity $v_{\mathrm{kick}}$ is a vector that plays an important if not the decisive role in the Gaia BH-like systems formation. What counts is not only the magnitude of the kick but also its direction and relative position of the stars on their mutual orbit when SN takes place.
@@
The natal kick of a given magnitude determines $v_{\mathrm{sys}}$ but not the eccentricity of the binary, what is demonstrated in Fig. 6. However, there is a correlation between the eccentricity, post- and pre-SN orbital separation (see e.g. Kalogera, 1996) which is followed by our population of Gaia BH-like binaries as is shown in Fig. 7. For Gaia BH1, where the companion star is still on its main sequence, the evolution has not have enough time to change the binary orbit from its immediate post-SN configuration. The system may have widen or shrunken due to the natal kick but the total amount of the orbital separation change is tightly connected to the system post-SN eccentricity. If the natal kick causes binary orbit to widen the orbital separation may increase no more than twice ($a_{\mathrm{post-SN}}/a_{\mathrm{pre-SN}}\sim 2$) for binaries that have eccentricities $e\in[0.4;0.6]$. If the orbit shrinks due to the natal kick then it can shrink by no more than $40\%$ ($1.00.6<a_{\mathrm{post-SN}}/a_{\mathrm{pre-SN}}<$ 1.0) (see Fig. 7) for such binaries.
@@
Whether the binary orbit shrinks or expands due to the natal kick depends on the position of the exploding star on the orbit around its companion at the moment of the SN explosion and on the natal kick vector, where both the position and the natal kick vector are drawn from the distributions described in Sec. 2.1.
Apart from the shrinkage/widening of the orbit the natal kick changes the angle between the direction of BH spin and the direction of the binary angular momentum ($\theta$). We show the relation between angles $\theta$, $\gamma$ and $v_{\mathrm{sys}}$ (top) and $a_{\mathrm{pre-SN}}$/$a_{\mathrm{post-SN}}$ (bottom) for Gaia BH1-like and Gaia BH2-like binaries in Fig. 8. On both plots Gaia BH1-like binaries are marked with stars and Gaia BH2-like binaries are marked with filled circles. The formation of Gaia BH1-like binaries depends differently on $\gamma$ than the formation of Gaia BH2-like binaries.
@@
In the case of Gaia BH2-like binaries $\gamma$ is a major factor determining their population. Most of them ($\sim 95\%$) form if natal kick direction is close to the orbital plane ($\gamma\sim 90^{\circ}\pm 15$) and $v_{\mathrm{kick}}<41.6$ km/s. The binaries that survive SN and form Gaia BH2-like systems have $\theta$ in the whole range of values. It can be clearly seen from the $\gamma$ and $\theta$ distributions shown in Fig. 10: $\theta$ distribution is almost flat and $\gamma$ distribution is centered around $90^{\circ}$.
@@
The evolution Gaia BH-like binaries: the tidal interactions.
To calculate the evolution of the orbital separation and eccentricity due to tidal interaction between the binary components, the convective damping in equilibrium tide and radiative damping in the dynamical tide (Zahn, 1977) are taken into account following the formalism of Hut (1981). The discrepancies between the tidal circularization periods derived from observations of coeval binary stars populations (Meibom & Mathieu, 2005) and from synthetic binary populations have led to the introduction of an additional scaling factor, $F_{\mathrm{tid}}$, in the tidal evolution equations of orbital separation and eccentricity in StarTrack. It appeared that to match the observed cut off period in the open cluster M67 with simulations, the scaling factor has to be set to $F_{\mathrm{tid}}=50$ in the case of convective damping (Belczynski et al., 2008). Although different approaches has been proposed to determine the circularization periods of binary populations (e.g. Zanazzi 2022; Bashi et al. 2023) and there are also new approaches to the theory of tides (e.g. Terquem 2021; Terquem & Martin 2021), it seems that the need of enhancement of convective damping is maintained (Mirouh et al., 2023).
To evaluate the impact of $F_{\mathrm{tid}}$ on our results on Gaia BH-like binary formation, we tested three values of the scaling factor: $F_{\mathrm{tid}}=1.0$, $50.0$ and $100.0$. We find that setting $F_{\mathrm{tid}}=1.0$ (the standard assumption in most binary population synthesis calculations) reduces the number of Gaia BH-like binaries in the “V2 standard” model by a factor of $3.2$ (from $5550$ to $1757$ binaries), while $F_{\mathrm{tid}}=100$ increases that number only by $5\%$ (from $5550$ to $5815$). Changing the efficiency of tidal interaction in the convective case impacts mainly the binaries that we classify as Gaia BH1-like: their number drops from $219$ for $F_{\mathrm{tid}}=100.0$ to $51$ for $F_{\mathrm{tid}}=1.0$. However, it has no such effect on Gaia BH2-like binaries ($166$ binaries for $F_{\mathrm{tid}}=1.0$, $73$ for $F_{\mathrm{tid}}=50.0$ and $96$ for $F_{\mathrm{tid}}=100.0$). The reason for the reduction in the number of Gaia BH1-like binaries can be seen from their evolutionary path. At first, the orbit of the system widens due to the wind mass-loss from the massive primary star and, after a few Myr, the star evolves into the red supergiant phase when the radius becomes larger than $1000\,\mathrm{R_{\odot}}$ and the star’s envelope becomes convective. This is when the tidal interaction between the stars sets in and the efficiency of the convective damping impacts the subsequent binary evolution. For $F_{\mathrm{tid}}=50.0$, the efficient convective damping causes the orbital separation to start shrinking and, at the point when the evolved star’s radius reaches almost $2000\,\mathrm{R_{\odot}}$, the binary components are close enough to enter a CE phase. On the other hand, for $F_{\mathrm{tid}}=1.0$, the orbit does not shrink at all. The binary widens throughout the evolution of the massive primary due to the latter’s stellar wind, until supernova explosion disrupts the binary. That way, the initial parameter space for the binaries to form Gaia BH1-like system is much more limited for $F_{\mathrm{tid}}=1.0$.
There is no such effect in the case of Gaia BH2-like binaries because they are much wider than Gaia BH1-like binaries and most of them do not need to get close enough to evolve through a CE phase. The formation of Gaia BH2-like binaries is mainly determined by the magnitude and the direction of the natal kick. The difference between the $F_{\mathrm{tid}}=1.0$ and the $F_{\mathrm{tid}}=100.0$ models is that most of the progenitors of Gaia BH2-like binaries start their evolution with smaller orbital separations ($a_{0}=4000-7000\,\mathrm{R_{\odot}}$) in the former case than in the latter case ($a_{0}=6000-9000\,\mathrm{R_{\odot}}$).
The evolution Gaia BH-like binaries: the common envelope parameters
The results from the IBE evolution channel depend on the choice of two parameters that determine the CE evolution of the binary, namely, the envelope binding energy parameter $\lambda_{\mathrm{ce}}$ and the envelope efficiency parameter $\alpha_{\mathrm{ce}}$. In this section, we discuss the choice of $\lambda_{\mathrm{ce}}$ and $\alpha_{\mathrm{ce}}$ for our models.
As was mentioned in Sec.2.1, for the calculations done for the purpose of this paper we adopt the $\lambda_{\mathrm{ce}}$ formulae by Wang et al. (2016) for stars more massive than $18.0\,\mathrm{M_{\odot}}$, which we find to be more complete compared to the older prescriptions.The reason is that to calculate the binding energy of an envelope, Wang et al. (2016) take into account not only the contributions from the gravitational binding energy, internal energy of the envelope consisting of the thermal energy, radiation energy, and ionization and dissociation energies but also the enthalpy (see Ivanova & Chaichenets, 2011, and discussion therein). The inclusion of enthalpy decreases the envelope’s binding energy with respect to the cases where it is not taken into account. This approach differs from the commonly used $\lambda_{\mathrm{ce}}$ prescription given by Claeys et al. (2014), where only gravitation and ionization are considered as the energy sources that unbind the donor’s envelope. To see how different choices of $\lambda_{\mathrm{ce}}$ impact the envelope binding energy for stars of different masses and metallicities, we refer to Fig. A1 in the Appendix of Iorio et al. (2023), although it does not present the case of the Wang et al. (2016) formulae. From that figure, it can be seen that, for a $M_{\mathrm{ZAMS}}\sim 30\,\mathrm{M_{\odot}}$ star of metallicity $Z=0.01$, the envelope binding energy during the core He-burning phase is $\sim 2$ orders of magnitude smaller for the $\lambda_{\mathrm{ce}}$ of Claeys et al. (2014) than for the $\lambda_{\mathrm{ce}}$ of Xu & Li (2010b). Consequently, the envelope ejection during a CE phase is easier in the former than in the latter case, for the same value of $\alpha_{\mathrm{ce}}$ in both cases. This, in turn, may result in different post-CE orbital separations in models where, apart from $\lambda_{\mathrm{ce}}$, all other assumptions are identical.
@@
The standard version of StarTrack has been using $\lambda_{\mathrm{ce}}$ formulae from Xu & Li (2010a); Xu & Li (2010b). However, for the stars with ZAMS masses in the range of $\sim 35-45\,\mathrm{M_{\odot}}$, which are the progenitors of $\sim 8-10\,\mathrm{M_{\odot}}$ BHs, those formulae give $\lambda_{\mathrm{ce}}$ values as high as $4.0$ for the core He burning stars because they underestimate their envelope binding energy. The reason is that the formulae in Xu & Li (2010a); Xu & Li (2010b) were fitted to the models calculated with the old Eddington’s code EV only for stars with $M_{\mathrm{ZAMS}}<20\,\,\mathrm{M_{\odot}}$. Those formulae were updated using Modules for Experiments in Stellar Astrophysics code (MESA, Paxton et al. (2011)) in 2016 by Wang et al. (2016) and should be used in their most recent form (Xiang-Dong Li private communication). We further motivate our choice of the Wang et al. (2016) $\lambda_{\mathrm{ce}}$ by the fact that the formulae were fitted using a dense grid of stellar models calculated with MESA. Furthermore, they assumed a self-consistent core-envelope boundary located in the hydrogen-burning shell of the maximum compression prior to the CE, following Ivanova (2011) (see, however, Klencki et al., 2021). The formulae from Wang et al. (2016), that we use, were calculated using the prescription for the wind mass loss rates from Hurley et al. (2000) and Vink et al. (2001) (Wind1 in Wang et al. (2016)) and we take $\lambda_{\mathrm{ce}}$ as the average of $\lambda_{\mathrm{g}}$ (corresponding to gravitational binding energy only), $\lambda_{\mathrm{b}}$ (corresponding to the total energy, comprising both gravitational and internal energy) and $\lambda_{\mathrm{h}}$ (which, in addition to the total energy, takes into account the enthalpy of the stellar envelope).
@@
Note that our results are obtained with the standard value of $\alpha_{\mathrm{ce}}$, i.e., $\alpha_{\mathrm{ce}}=1.0$. There are analytical estimates for the final orbital separation, after the CE phase in pre-Gaia BH1 systems (e.g. El-Badry et al., 2023a). However, they are based on the assumption that the donor star has to lose a mass of $20\,\mathrm{M_{\odot}}$ in envelope, during the CE. This is an overestimation, considering that the BH progenitor loses up to half of its mass in stellar winds during its evolution, before the CE phase starts (see Fig. 1 and its description in Sec.2.2.2). Consequently, the envelope mass, that has to be unbound during the CE phase, has a mass of only $\sim 3\,\mathrm{M_{\odot}}$. If $M_{\mathrm{env}}=3.0\,\mathrm{M_{\odot}}$ is considered in the calculations instead, the post-CE orbital separation of the binary turns out to be $>100\,\mathrm{R_{\odot}}$ instead of a few $\mathrm{R_{\odot}}$.
Taking into account the uncertainties of the $\lambda_{\mathrm{ce}}$ prescription and the correction to the post-CE orbital separation calculations as mentioned above, we conclude that invoking $\alpha_{\mathrm{ce}}>1$ should not be necessary to form Gaia BH-like binaries via the IBE channel. Also, it is so far unclear what (if any) physical mechanism could lead to $\alpha_{\mathrm{ce}}>1$. Therefore, we assume $\alpha_{\mathrm{ce}}=1$, which still seems to be an optimistic assumption as it implies that all orbital energy goes into unbinding the envelope without any energy loss.
For the analyses of the impact of different $\alpha_{\mathrm{ce}}$ values, of different SN engines (e.g., rapid and delayed), and of the reduction of SN natal kicks due to matter fallback on the isolated-binary formation of BH–star binaries, we refer to the works of Yamaguchi et al. (2018) and Chawla et al. (2022).
@@ 2.3.4 Gaia BH1 as Low Mass X-ray Binary
There is no X-ray emission detected from Gaia BH1 and Gaia BH2. However, it is found that during the post-dormant BH binary phase, both of these binaries may evolve through the symbiotic X-ray binary phase to the phase of stable Roche-lobe overflow (RLOF) Rodriguez et al. (2023). At this point, the systems convert into low mass X-ray binaries (LMXBs). The population of LMXBs, that could arise from Gaia BH-like systems, would have long orbital periods of $P_{\mathrm{orb}}>100$ days.
@@ 3.2 Results: formation rate of field BH-star binaries
As evident from Figs. 11 and 12, the vast majority of dynamically assembled field BH-star systems have $M_{\rm comp}<3\mathrm{M_{\odot}}$. Nearly half of them have zero or small eccentricity, an MS companion, and $P_{\rm orb}<1$ day — these very tight binaries are outcomes of tidal capture between the BH and the companion star inside the parent cluster and the binary’s subsequent tidal circularization [^3] In the Ba21 runs, tidal energy loss and analytical tidal circularization were enabled (option $27=1$; Aarseth 2003).. Inside clusters, such close binaries behave like single, massive objects most of the time and are ejected dynamically in ways similar to those of ejection of single BHs (Kulkarni et al., 1993). These systems are potentially tidally interacting or mass transferring and are not compatible with the Gaia detected, detached BH-star binaries. There are also a few wide, $P_{\rm orb}\gtrsim 100$ day, MS-companion binaries — these systems are assembled in the cluster via dynamical exchange encounters and are also ejected via dynamical scatterings (see above), making them eccentric in general.
@@
Since the model clusters produce only one Gaia BH1- and one Gaia BH2-like binary (see above), we do not repeat the source count analysis in Sec. 2.3.2 for the cluster channel. We note that the relative contribution of star clusters to the field Gaia-BH binaries depends on cluster formation efficiency (hereafter CFE), i.e., the relative fractions of the star forming mass that go into star clusters and isolated binaries. An appropriate CFE for a Milky Way-like spiral galaxy is largely an open question (see, e.g., Krumholz et al. 2019 for a review). The recent study by Niccolò Di Carlo et al. (2023) adopted a CFE of 10%. Notably, the number, $\sim 900$, of isolated-binary Gaia BH systems, as obtained in Sec. 2.3.2, corresponds to the highest formation efficiency from the isolated-binary channel, which formation efficiency is similar to that for the cluster channel (c.f. Tables 3 and 5). Hence, the 5-year count of Gaia BH-like systems from clusters would simply proportionate with the CFE, with respect to the above isolated-binary-channel count. A detailed study of CFE in the appropriate cluster mass range is beyond the scope of this work.
@@ 4 Conclusions
In this paper we have investigated the two possible channels of the formation of dormant BH binaries Gaia BH1 and Gaia BH2: the isolated binary evolution (Sec. 2) and the dynamical interactions in young massive and open star clusters (Sec. 3). We used the population synthesis code StarTrack for IBE calculations (Sec. 2.1) and the updated star-by-star N-body integration code NBODY7 for cluster models (Sec. 3.1). The two systems from IBE Gaia-BH like binaries population, that match well the parameters of Gaia BH1 and Gaia BH2, were evolved till the Hubble time to find the characteristics of the low-mass X-ray binaries that possibly form from such systems (Sec.2.3.4). We have also analysed what are the constrains on magnitude and direction of the natal kick imparted to BH set by the orbital parameters and systemic velocities of Gaia BH1-like and Gaia BH2-like binaries (Sec. 2.2.3 and Sec. 2.3.3).
@@
The formation rate of Gaia BH-like binaries, as inferred from the present simulations, is sensitive to the criteria adopted to classify a binary as Gaia BH-like (Sec. 2.2.1) and it can differ by three orders of magnitude for the IBE channel (see Table 3). Ensuring that the common criteria are set and the same total stellar mass per simulation is adopted for IBE and dynamical interactions in young clusters, we find that the formation rate of Gaia BH-like binaries for both channels is equal $2.7\times 10^{-7}\,\mathrm{M_{\odot}}^{-1}$. If we consider the general population of field BH-low mass star ($<3\,\mathrm{M_{\odot}}$) binaries from young massive and open clusters and corresponding BH-low mass star population from IBE the formation rate of such binaries in both channels is on order of $\sim 10^{-6}\,\mathrm{M_{\odot}}^{-1}$.
Our approximate estimate is that $\sim 900$ dormant BH binaries, formed through the IBE channel, would be detectable by Gaia in the Milky Way thin disc.
@@
  • The population of Gaia BH1-like binaries is not vulnerable to the natal kick velocity distribution adopted but a seven fold increase in the number of Gaia BH2-like binaries in model V2 compared to model V1 favours the two-component Maxwellian distribution with low- and high-velocity components ($\sigma\approx 21$ km/s and $\sigma\approx 107$ km/s), as suggested by Zhao et al. (2023).
  • The median value of natal kick velocity imparted to the systems, that become Gaia BH1-like and Gaia BH2-like binaries, is $\sim 39$ km/s for the former and $\sim 19$ km/s for the latter (assuming the natal kick distribution from Zhao et al. (2023)).
@@
  • For Gaia BH1-like binary to be formed, the kicks should be directed out of the binary orbital plane: the higher is the natal kick velocity the further from orbital plane should the kick be directed.
@@
We find that, in addition to the natal kicks, the treatment of the tidal interactions in binaries should not be overlooked in the context of the formation of Gaia BH-like binaries. The enhancement of the convective damping efficiency relative to the classical prescription Hut (1981), introduced as the additional factor $F_{\mathrm{tid}}$ in the equations, increases the number of the systems like Gaia-BH1 in the IBE channel. However, increasing $F_{\mathrm{tid}}$ above the certain value, which was originally calibrated to match the observations ($F_{\mathrm{tid}}=50$, Belczynski et al., 2008), does not lead to further significant increase in the Gaia BH1-like binaries number (see Sec. 2.3.3). The results show that the prescriptions for tidal interactions that are used in the codes play a decisive role in the formation of Gaia BH1-like binaries via the IBE channel and that more detailed investigation on the subject is needed.